# Optimal Conformal Prediction, E-values, Fuzzy Prediction Sets and Subsequent Decisions

### v1

@@ 3.4 Example: fuzzy Gaussian prediction
Figure 3: Optimal fuzzy confidence sets for Neyman–Pearson-utility (solid), log-utility (dotted) and bounded log-utility (dashed) under the simple Gaussian setting (left) and composite Gaussian setting (right) from Example 1, 2 and 4, for $n=3$, $\mu=0$, $\sigma=1$, $\tau=3.5$, $\bar{Z}_{n}=1.44$. Here, the Neyman–Pearson utility and bounded log-utility are both bounded at $1/0.05=20$.
@@
In Section 2.4, we implicitly considered optimal fuzzy confidence sets under the classical Neyman–Pearson utility $x\mapsto x\wedge 1/\alpha$. We continue these examples under more general utility functions.
We first consider the log-utility version of Example 1, where we had the simple model $\mathcal{P}=\{\mathcal{N}(\mu 1_{n+1},\sigma^{2}I_{n+1})\}$ and optimized against the alternative $Q=\mathcal{N}(\mu 1_{n},\sigma^{2}I_{n})\times\mathcal{N}(\mu,\tau^{2})$, for known $\mu\in\mathbb{R}$ and $\tau>\sigma>0$. Here, we could choose $\tau\to\infty$ to mimic the Lebesgue measure, as in Section 2.4, but we find that this generally yields undesirable fuzzy confidence sets.
@@
$\varepsilon(z^{n+1})=\frac{d\mathcal{N}(\mu,\tau^{2})}{d\mathcal{N}(\mu,\sigma^{2})}(z_{n+1})=\mathcal{E}^{z^{n}}(z_{n+1}).$ (12)
For $1/\alpha$-bounded log-utility $U:x\mapsto\log(\varepsilon\wedge 1/\alpha)$, the optimal e-value is the likelihood ratio capped at $1/\alpha$ and ‘boosted’ by some constant $b_{\alpha}$ so that it has expectation exactly 1 (Koning, 2025b):
@@
In the left panel of Figure 3, we display these fuzzy confidence sets next to the traditional confidence set from Example 1.
@@
where $A=Z^{n}-\bar{Z}_{n}$ and $B=Z_{n+1}-\bar{Z}_{n}$. Fixing $\bar{Z}_{n}$ and evaluating this likelihood ratio at plug-in values $z_{n+1}\in\mathbb{R}$ for $Z_{n+1}$, we recover the log-optimal fuzzy confidence set:
@@
which relies on the fact that the Hunt-Stein theorem goes through for log-optimal e-values as proven by Pérez-Ortiz et al. (2022). To the best of our knowledge, this has not been proven yet for general expected-utility-optimal e-values, but we would be surprised if it does not go through in general.
Comparing the fuzzy confidence set in (13) to (12), we see that this likelihood ratio is centered at $\bar{Z}_{n}$ instead of $\mu$ and the variance in both numerator and denominator is inflated by an additional $\tau^{2}/n$- and $\sigma^{2}/n$-term due to the estimation of $\mu$ by $\bar{Z}_{n}$. We illustrate this fuzzy confidence set in the right panel of Figure 3, along its bounded version and the Neyman–Pearson-utility variant from Example 2.
@@ 4 Optimal conformal prediction
In this section, we specialize our methodology to conformal prediction, where $\mathcal{P}$ is the class of exchangeable distributions on $\mathcal{Z}^{n+1}$. We cover both optimality for classical conformal confidence sets and more general fuzzy conformal confidence sets.
As we showed in Section 3.2, a fuzzy conformal confidence set is equivalent to an e-value $\varepsilon$, which is valid for $\mathcal{P}$ when $\varepsilon$ is a valid e-value for $\mathcal{P}$. Moreover, in Section 3.3 we show that the fuzzy confidence set is optimal if $\varepsilon$ is optimal for some expected-utility target $\mathrm{E}^{Q}[U(\varepsilon)]$. As a consequence, we may reduce the problem of studying optimal (fuzzy) conformal prediction sets to studying optimal e-values for exchangeability. Optimal e-values for exchangeability (and more general group invariance) have recently been developed in Koning (2025a).
@@ 4.1 Background: orbits & exchangeability
Note that this is actually a statement about its distribution $P$, which we say is exchangeable if it is the law of an exchangeable random variable. We use $\mathcal{P}^{\textnormal{exch.}}$ to denote the collection of all exchangeable distributions on $\mathcal{Z}^{n+1}$.
@@
$P$ is exchangeable if and only if $P^{|O}=\textnormal{Unif}(O)$, for every $O\in\mathcal{O}^{n+1}$.
4.2 Reduction to hypotheses on orbits
The key idea in Koning (2025a) to obtain an optimal e-value for exchangeability is to partition the problem into a simpler problem on each orbit $O\in\mathcal{O}^{n+1}$. For this purpose, we consider a conditional e-value $\varepsilon_{O}:O\to[0,\infty]$ on each orbit $O\in\mathcal{O}^{n+1}$, defined as the restriction of our e-value $\varepsilon$ to $O$:
$\varepsilon_{O}(z^{n+1})=\varepsilon(z^{n+1}),\textnormal{ for every }z^{n+1}\in O.$
Passing to the collection of these conditional e-values is lossless, as the test $\varepsilon$ may be recovered from the conditional tests, by evaluating the conditional e-value at the data orbit $O(Z^{n+1})$:
$\varepsilon(Z^{n+1})=\varepsilon_{O(Z^{n+1})}(Z^{n+1}).$
The motivation for shattering our e-value into a collection of conditional e-values is Theorem 6, specialized from Theorem 5 in Koning (2025a). It shows that if we construct an optimal conditional e-value $\varepsilon_{O}$ on each orbit $O\in\mathcal{O}^{n+1}$, then the corresponding unconditional e-value $\varepsilon$ is optimal.
Theorem 6 (Koning (2025a)).
If $\varepsilon_{O}$ is valid for $\textnormal{Unif}(O)$ and optimal against $\mathrm{E}^{Q^{|O}}[U(\varepsilon)]$, for every $O$, then $\varepsilon$ is valid for $\mathcal{P}^{exch.}$ and optimal against $\mathrm{E}^{Q}[U(\varepsilon)]$. Moreover, this e-value $\varepsilon$ is optimal uniformly in mixtures over $Q^{|O}$.
4.3 Further reduction of hypotheses
By Theorem 6, it remains to derive an (orbit-conditional) optimal e-value for $\textnormal{Unif}(O)$ against $Q^{|O}$. In this section, we use the available structure of $Q$ to reduce the hypothesis even further.
Recall that we intend to optimize against an alternative of the form $Q=P_{*}^{Z^{n}}\otimes Q^{Z_{n+1}\mid Z^{n}}$, where $P_{*}^{Z^{n}}$ is the ‘true’ law of $Z^{n}$. While we do not know this law of $Z^{n}$, the fact that the true law of $Z^{n+1}$ is exchangeable implies $Z^{n}$ is exchangeable. This allows us to pin $Q^{|O}$ up to the marginal distribution $Q^{Z_{n+1}|O}$ of $Z_{n+1}$, as captured in Proposition 1. Its proof is given in Appendix A.2.
In words, Proposition 1 states that a draw from $Q^{|O}$ is equivalent to first sampling $Z_{n+1}$ using some distribution over the elements of the tuple $Z^{n+1}]$, and subsequently sampling $Z^{n}$ from the remainder uniformly without replacement.
@@ 4.3 Further reduction of hypotheses
$Q^{Z^{n}|O,Z^{n+1}}=\textnormal{UnifWR}([Z^{n+1}]\setminus\{Z_{n+1}\}),$
As a consequence of Proposition 1, the conditional likelihood ratio $\textnormal{LR}^{|O}$ given $O$ reduces to
$\textnormal{LR}^{|O}(z^{n+1})=\frac{dQ^{|O}}{d\textnormal{Unif}(O)}(z^{n+1})$ (14)
$=\frac{d(Q^{Z_{n+1}\mid O}\otimes\textnormal{UnifWR}([Z^{n+1}]\setminus\{Z_{n+1}\}))}{d(\textnormal{Unif}[Z^{n+1}]\otimes\textnormal{UnifWR}([Z^{n+1}]\setminus\{Z_{n+1}\}))}(z^{n+1})$ (15)
$=\frac{dQ^{Z_{n+1}\mid O}}{d\textnormal{Unif}[Z^{n+1}]}(z_{n+1})=(n+1)\times Q^{Z_{n+1}\mid O}(z_{n+1})$ (16)
$=:\textnormal{LR}^{Z_{n+1}|O}(z_{n+1}),$ (17)
where $\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}:[z^{n+1}]\to\mathbb{R}_{+}$ is a function only of the final element $z_{n+1}$ of the tuple $z^{n+1}$.
To reduce the hypotheses, we may use the finding of Koning (2025b) that expected-utility-optimal e-values are non-decreasing functions of likelihood ratios. As the likelihood ratio reduces to a likelihood ratio of $\textnormal{Unif}[Z^{n+1}]$ to $Q^{Z_{n+1}\mid O}$, this means the optimal problem implicitly reduces to testing the hypotheses
$H_{0}^{O}$ $:Z_{n+1}\sim\textnormal{Unif}[Z^{n+1}],$
$H_{1}^{O}$ $:Z_{n+1}\sim Q^{Z_{n+1}\mid O},$
where $\textnormal{Unif}[Z^{n+1}]$ is well-defined because the orbit representative $Z^{n+1}]$ is uniquely identified by the orbit $O$.
The overall reduction of the alternative from an arbitrary probability measure $Q$ on $\mathcal{Z}^{n+1}$ to $Q^{Z_{n+1}\mid O}$ greatly reduces the complexity of the object we must specify.
@@
As optimal e-values only depend on the position of $Z_{n+1}$ on $Z^{n+1}]$, we may equivalently express all the results in terms of the rank of $Z_{n+1}$ among $Z^{n+1}]$.
To define a rank for arbitrary data, we may define it relative to the orbit representative $Z^{n+1}]$. In particular, we define $\textnormal{Rank}(Z^{n+1})$ as the index at which each element in the tuple $Z^{n+1}$ needs to be placed to obtain the representative $Z^{n+1}]$ of $O$. This specializes to the conventional rank for data that has a natural order, if $Z^{n+1}]$ is sorted in ascending order.
Let $R_{n+1}^{O}$ denote last element of the tuple $\textnormal{Rank}(Z^{n+1})$, which corresponds to the rank of $Z_{n+1}$ among $Z^{n+1}]$. We may then equivalently express the reduced hypotheses as
@@
4.4 Independence under alternative
In Proposition 2 we further specialize from an exchangeable alternative to an alternative with (unconditional) independence. Concretely, the result implies that if we have an alternative of the form $Q=(P_{*}^{Z^{1}})^{n}\times\mu$, then
$\textnormal{LR}^{|O(z^{n+1})}(z^{n+1})\propto\frac{d\mu}{dP_{*}^{Z^{1}}}(z_{n+1}).$
If we interpret $\mu$ as some reference measure, a consequence of this result is that optimal e-values will be decreasing in the inverse of the ‘true’ $P_{*}^{Z^{1}}$-density of $z_{n+1}$. We believe this makes intuitive sense, as values of $z_{n+1}$ that have low density may be interpreted as being ‘outlying’ or ‘non-conforming’.
@@ 4.4 Independence under alternative
Suppose that $Q$ is so that $Y^{n+1}$ is independent and $Y^{n}$ is i.i.d., with $Q^{Z_{n+1}}$ and $Q^{Z_{1}}$ mutually absolutely continuous. Then,
$\textnormal{LR}^{|O(z^{n+1})}(z^{n+1})$
$=\frac{dQ^{Z_{n+1}}}{dQ^{Z_{1}}}(z_{n+1})\Bigg{/}\left(\frac{1}{(n+1)}\sum_{i=1}^{n+1}\frac{dQ^{Z_{n+1}}}{dQ^{Z_{1}}}(z_{i})\right).$
4.5 Optimal fuzzy conformal
Now that we have reduced the alternative, we may derive expected-utility-optimal e-values for conformal prediction.
Under some regularity conditions on the utility function $U$, we may obtain a clean and interpretable result, which we present in Theorem 7. Note that the conditional likelihood ratio is central in this optimality result. The result follows our reduction to the hypotheses $H_{0}^{O}$ and $H_{1}^{O}$, and a combination of Theorem 6 with Theorem 1 in Koning (2025b) on the shape of expected-utility-optimal e-values.
Theorem 7.
Assume $U$ is concave, non-decreasing, differentiable, $U(0)=0$, with continuous and strictly decreasing derivative $U^{\prime}$, which has limits $U^{\prime}(0)=\infty$ and $U^{\prime}(\infty)=0$. Then, an expected utility-optimal e-value is given by
$\varepsilon^{U}(z^{n+1})=(U^{\prime})^{-1}\left(\lambda^{O(z^{n+1})}\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}(z_{n+1})\right),$
where $\lambda^{O}\geq 0$ is an orbit-dependent normalization constant. This e-value is expected-utility-optimal, uniformly in any mixture over orbits $O$ and any true exchangeable distribution of $Z_{n}$.
Example 5 (Generalized-mean-optimality).
An example of a utility function that has been primarily studied in the e-value literature is $U:x\mapsto\log(x)$, which yields the conditional likelihood ratio itself as the optimal e-value:
$\varepsilon^{\log}(z^{n+1})=\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}(z_{n+1}).$
This generalizes to power-utility $U:x\mapsto(x^{h}-1)/h$, $h<1$, where the optimizer equals
$\varepsilon^{h}(Y_{n+1})=\frac{(\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}(z_{n+1}))^{\frac{1}{1-h}}}{\frac{1}{n+1}\sum_{j=1}^{n+1}(\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}(z_{n+1}))^{\frac{1}{1-h}}}.$
The parameter $h$ can be interpreted as tuning how ‘risky’ the e-value is, where large values of $h$ (close to 1) yield a risky e-value and small values ($h\to$-$\infty$) a less risky e-value. Indeed, assuming $\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}>0$, then the $h\to$-$\infty$ limit is the no-risk e-value
$\lim_{h\to-\infty}\varepsilon^{h}\equiv 1.$
Moreover, assuming $\textnormal{LR}^{Z_{n+1}|O}$ is uniquely maximized at $z_{*}^{O}$, the $h\to 1$ limit is the most risky e-value that goes all-in on a single element:
$\lim_{h\to 1}\varepsilon^{h}(z^{n+1})=\begin{cases}n+1,&\textnormal{ if }z_{n+1}=z_{*}^{O(z^{n+1})},\\ 0,&\textnormal{ otherwise}.\end{cases}$
Example 6 (Clipping and dampening e-values).
$\varepsilon(z^{n+1})=\left(\lambda\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}(z_{n+1})\right)\vee b,$
where $\lambda$ is some normalization constant that ensures $\varepsilon$ is an exact e-value. This may be interpreted as ‘clipping’ the likelihood ratio from below, and simultaneously shrinking it by $\lambda$ to ensure it remains a valid e-value. This may be proven analogously to the proof of upper-bounded e-values covered in Koning (2025b).
An alternative strategy to ‘dampen’ e-values is introduced by Grünwald (2023b). There, he proposes to take an e-value $\varepsilon$ and dampen it to the e-value $b+(1-b)\varepsilon$ that may be interpreted as a $b$-mixture of the e-value and the constant 1. If $\varepsilon$ is the log-utility maximizer, then we find $b+(1-b)\varepsilon$ implicitly maximizes the utility function
$U(\varepsilon)=\log([\varepsilon-b]\vee 0).$

### v2

@@ 4.5 Example: fuzzy Gaussian prediction
Figure 4: Optimal fuzzy prediction sets for Neyman–Pearson-utility (solid), log-utility (dotted) and bounded log-utility (dashed) under the simple Gaussian setting (left) and composite Gaussian setting (right) from Example 1, 2 and 5, for $n=3$, $\mu=0$, $\sigma=1$, $\tau=3.5$, $\bar{Z}_{n}=1.44$. Here, the Neyman–Pearson utility and bounded log-utility are both bounded at $1/0.05=20$.
@@
In Section 3.5, we implicitly considered optimal fuzzy prediction sets under the classical Neyman–Pearson utility $x\mapsto x\wedge 1/\alpha$. We continue these examples under more general utility functions.
We first consider the log-utility version of Example 1, where we had the simple model $\mathcal{P}=\{\mathcal{N}(\mu 1_{n+1},\sigma^{2}I_{n+1})\}$ and optimized against the alternative $Q=\mathcal{N}(\mu 1_{n},\sigma^{2}I_{n})\times\mathcal{N}(\mu,\tau^{2})$, for known $\mu\in\mathbb{R}$ and $\tau>\sigma>0$. Here, we could choose $\tau\to\infty$ to mimic the Lebesgue measure, as in Section 3.5, but we find that this generally yields undesirable fuzzy prediction sets.
@@
$\mathcal{E}(z^{n+1})=\frac{d\mathcal{N}(\mu,\tau^{2})}{d\mathcal{N}(\mu,\sigma^{2})}(z_{n+1})=\mathcal{E}^{z^{n}}(z_{n+1}).$ (12)
For $1/\alpha$-bounded log-utility $U:x\mapsto\log(\mathcal{E}\wedge 1/\alpha)$, the optimal e-value is the likelihood ratio capped at $1/\alpha$ and ‘boosted’ by some constant $b_{\alpha}$ so that it has expectation exactly 1 (koning2025continuoustestingunifyingtests):
@@
In the left panel of Figure 4, we display these fuzzy prediction sets next to the traditional prediction set from Example 1.
@@
where $A=Z^{n}-\bar{Z}_{n}$ and $B=Z_{n+1}-\bar{Z}_{n}$. Fixing $\bar{Z}_{n}$ and evaluating this likelihood ratio at plug-in values $z_{n+1}\in\mathbb{R}$ for $Z_{n+1}$, we recover the log-optimal fuzzy prediction set:
@@
which relies on a variant of the Hunt-Stein theorem derived for log-optimal e-values proven by perez2022statistics. To the best of our knowledge, this has not been proven yet for general expected-utility-optimal e-values, but we would be surprised if it does not go through in general.
Comparing the fuzzy prediction set in (13) to (12), we see that this likelihood ratio is centered at $\bar{Z}_{n}$ instead of $\mu$ and the variance in both numerator and denominator is inflated by an additional $\tau^{2}/n$- and $\sigma^{2}/n$-term due to the estimation of $\mu$ by $\bar{Z}_{n}$. We illustrate this fuzzy prediction set in the right panel of Figure 4, along its bounded version and the Neyman–Pearson-utility variant from Example 2.
@@ 5 Optimal conformal prediction
In this section, we specialize our methodology to conformal prediction, where $\mathcal{P}$ is the class of exchangeable distributions on $\mathcal{Z}^{n+1}$. We cover both optimality for classical conformal prediction sets and more general fuzzy conformal prediction sets.
As we showed in Section 4.2, a valid fuzzy conformal prediction set for the model $\mathcal{P}$ corresponds to a valid e-value $\mathcal{E}$ for ‘hypothesis’ $\mathcal{P}$. Moreover, in Section 4.3 we show that the fuzzy prediction set is optimal if $\mathcal{E}$ is optimal for some expected-utility target $\mathrm{E}^{Q}[U(\mathcal{E})]$. As a consequence, we may reduce the problem of studying optimal (fuzzy) conformal prediction sets to studying optimal e-values for exchangeability. Optimal e-values for exchangeability (and more general group invariance) have recently been developed in koning2023post.
@@ 5.1 Background: orbits & exchangeability
Note that this is actually a statement about its distribution $P^{Z^{n+1}}$, which we say is exchangeable if it is the law of an exchangeable random variable. We use $\mathcal{P}^{\textnormal{exch.}}$ to denote the collection of all exchangeable distributions on $\mathcal{Z}^{n+1}$.
@@
$P$ is exchangeable if and only if $P^{|O}=\textnormal{Unif}(O)$, for $P^{O}$-almost every orbit $O\in\mathcal{O}^{n+1}$.
5.2 Optimal fuzzy conformal
The key idea in koning2023post to obtain an optimal e-value for exchangeability is to simplify the problem to finding a ‘locally’ optimal e-value $\mathcal{E}_{O}:O\to[0,\infty]$ on each orbit $O\in\mathcal{O}^{n+1}$, and subsequently recombining these ‘local’ e-values into a ‘globally’ optimal e-value:
$\mathcal{E}(Z^{n+1})=\mathcal{E}_{O(Z^{n+1})}(Z^{n+1}).$
Indeed, Theorem 6, which is specialized from Theorem 5 in koning2023post, shows that if each local e-value is optimal, then the resulting unconditional e-value $\mathcal{E}$ is optimal.
Theorem 6 (koning2023post).
If $\mathcal{E}_{O}$ is valid for $\textnormal{Unif}(O)$ and optimal for $\mathrm{E}^{Q^{|O}}[U(\cdot)]$, for every $O$, then $\mathcal{E}$ is valid for $\mathcal{P}^{exch.}$ and optimal against $\mathrm{E}^{Q}[U(\mathcal{E})]$. Moreover, this e-value $\mathcal{E}$ is optimal uniformly in mixtures over $Q^{|O}$.
Theorem 6 reduces the problem to constructing a local e-value $\mathcal{E}_{O}$ for $\textnormal{Unif}(O)$ against $Q^{|O}$. To find such a local e-value, we use the “Neyman–Pearson lemma for e-values” recently derived in koning2025continuoustestingunifyingtests. The classical Neyman–Pearson lemma corresponds to the choice $U(x)=x\wedge 1/\alpha$. The version we present in Theorem 7 specializes to the structure at hand: $O$ is finite and $Q^{|O}$ is absolutely continuous with respect to $\textnormal{Unif}(O)$, since $\textnormal{Unif}(O)$ has full support. Corollary 1 presents a more easily interpretable version under some regularity conditions on $U$ and $Q^{|O}$.
Theorem 7 (koning2025continuoustestingunifyingtests).
Let $U:[0,\infty]\to[-\infty,\infty]$ be concave, non-decreasing and upper-semicontinuous. Then, the optimization problem
$\sup_{\mathcal{E}}E^{Q^{|O}}[U(\mathcal{E})],\textnormal{ s.t. }\mathcal{E}:\mathcal{X}\to[0,\infty],\ E^{\textnormal{Unif}(O)}[\mathcal{E}]\leq 1,$
admits an optimal solution. Moreover, if $\mathcal{E}^{*}$ is an optimizer, then there exists a normalization constant $\lambda_{O}\geq 0$ such that
$\lambda_{O}\bigg/\frac{dQ^{|O}}{d\textnormal{Unif}(O)}$ $\in\partial U(\mathcal{E}^{*}),$
on $\{Q^{|O}>0\}$ and $\mathcal{E}^{*}=0$ on $\{Q^{|O}=0\}$, where $\partial U$ is the superdifferentiable of $U$.
Corollary 1.
If $Q^{|O}$ has full support on $O$ and $U$ is differentiable with strictly decreasing derivative $U^{\prime}$, then an optimizer is
$\mathcal{E}^{*}=(U^{\prime})^{-1}\left(\lambda_{O}\bigg/\frac{dQ^{|O}}{d\textnormal{Unif}(O)}\right),$
where $(U^{\prime})^{-1}$ denotes the functional inverse of $U^{\prime}$.
A consequence of Theorem 7 is that an optimal local e-value $\mathcal{E}_{O}$ only depends on the data through the orbit-conditional likelihood ratio $\textnormal{LR}^{|O}:=dQ^{|O}/d\textnormal{Unif}(O)$. In Proposition 1, we use the structure of $Q=P^{Z^{n}}\otimes\mu^{|Z^{n}}$ to show that this conditional likelihood ratio only depends on $Z^{n+1}$ through $Z_{n+1}$. Its proof may be found in Appendix A.2.
@@ 5.2 Optimal fuzzy conformal
$\textnormal{LR}^{|O}(z^{n+1})$ $=\frac{dQ^{|O}}{d\textnormal{Unif}(O)}(z^{n+1})$
$=\frac{Q^{Z_{n+1}|O}(z_{n+1})}{1/(n+1)}=:\textnormal{LR}^{Z_{n+1}|O}(z_{n+1}).$
A consequence of Proposition 1 is that an expected utility optimal local e-value $\mathcal{E}_{O}^{*}$ does not depend on $Z^{n}$. Indeed, we are implicitly considering the following hypotheses on each orbit:
$H_{0}^{O}$ $:Z_{n+1}\sim\textnormal{Unif}[Z^{n+1}]$
$H_{1}^{O}$ $:Z_{n+1}\sim Q^{Z_{n+1}|O}.$
By extension, a globally optimal e-value $\mathcal{E}^{*}$ only depends on $Z^{n}$ through the orbit $O(Z^{n+1})$.
@@
As optimal e-values only depend on the position of $Z_{n+1}$ in $Z^{n+1}]$, we may equivalently express all the results in terms of the rank of $Z_{n+1}$ among $Z^{n+1}]$.
To define a rank for arbitrary (non-numerical) data, we may define it relative to the orbit representative $Z^{n+1}]$. In particular, we define $\textnormal{Rank}(Z^{n+1})$ as the index at which each element in the tuple $Z^{n+1}$ needs to be placed to obtain the representative $Z^{n+1}]$ of $O$. This specializes to the conventional rank for data that has a natural order, if $Z^{n+1}]$ is sorted in ascending order.
Let $R_{n+1}^{O}$ denote last element of the tuple $\textnormal{Rank}(Z^{n+1})$, which corresponds to the rank of $Z_{n+1}$ among $Z^{n+1}]$. We may then equivalently express the reduced orbit-level hypotheses as
@@
5.3 Illustration: i.i.d. and split-conformal
As the conditional likelihood ratio is central in finding the optimal e-value, we first illustrate what such a conditional likelihood ratio may look like in a popular setting, before we proceed with several examples of utility functions.
In Proposition 2, we consider optimality against an i.i.d. alternative of the form $Q=P^{Z^{n}}\times\mu$ where $P^{Z^{n}}=(P^{Z_{1}})^{n}$. For example, such an alternative may arise by plugging-in an i.i.d. estimator $\widehat{P}^{Z^{n}}=(\widehat{P}^{Z_{1}})^{n}$ for $P_{*}^{Z^{n}}$, based on a separate sample akin to the popular ‘split-conformal prediction’ approach. Its proof is found in Appendix A.3.
@@
Suppose that $Q$ is so that $Z^{n+1}$ is independent and $Z^{n}$ is i.i.d., with $Q^{Z_{n+1}}$ and $Q^{Z_{1}}$ mutually absolutely continuous. Then,
$\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}(z_{n+1})$
$=\frac{dQ^{Z_{n+1}}}{dQ^{Z_{1}}}(z_{n+1})\Bigg/\left(\frac{1}{n+1}\sum_{i=1}^{n+1}\frac{dQ^{Z_{n+1}}}{dQ^{Z_{1}}}(z_{i})\right).$
Proposition 2 may be interpreted as stating that
$\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}(z_{n+1})\propto\frac{d\mu}{d\widehat{P}^{Z_{1}}}(z_{n+1}),$
where the proportionality is up to some orbit-dependent constant. Viewing $\mu$ as some reference measure, an interpretation of this result is that the optimal e-value is decreasing in the marginal $\widehat{P}^{Z_{1}}$-density at $z_{n+1}$. We believe this makes intuitive sense, as values of $z_{n+1}$ that have low density may be interpreted as being ‘outlying’ or ‘non-conforming’.
5.4 Generalized-mean utilities
In the e-value literature, the near-universal choice of utility is the log-utility $U:x\mapsto\log(x)$, for which the optimal e-value is the orbit-conditional likelihood ratio itself:
$\mathcal{E}^{\log}(z^{n+1})=\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}(z_{n+1}).$
This generalizes to the power-utility $U:x\mapsto(x^{h}-1)/h$, $h<1$, $h\neq 0$, where the optimizer equals
$\mathcal{E}^{h}(Y_{n+1})=\frac{(\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}(z_{n+1}))^{\frac{1}{1-h}}}{\frac{1}{n+1}\sum_{j=1}^{n+1}(\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}(z_{n+1}))^{\frac{1}{1-h}}}.$
The log-utility appears as the $h\to 0$ limit.
The parameter $h$ can be interpreted as tuning how ‘risky’ the e-value is, where large values of $h$ (close to 1) yield a risky e-value and small values ($h\to$-$\infty$) a less risky e-value. Indeed, assuming $\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}>0$, the $h\to$-$\infty$ limit is the no-risk e-value
$\lim_{h\to-\infty}\mathcal{E}^{h}\equiv 1.$
Moreover, assuming $\textnormal{LR}^{Z_{n+1}|O}$ is uniquely maximized at $z_{*}^{O}$, the $h\to 1$ limit is the most risky e-value that is large on a single element on the orbit and zero elsewhere:
$\lim_{h\to 1}\mathcal{E}^{h}(z^{n+1})=\begin{cases}n+1,&\textnormal{ if }z_{n+1}=z_{*}^{O(z^{n+1})},\\ 0,&\textnormal{ otherwise}.\end{cases}$